# Revisiting Power System Stabilizers with Increased Inverter-Based Generation: A Case Study

### v1

@@ Abstract
Index Terms:
grid-following converters, power system stabilizers, small-signal rotor angle stability, tuning methods.
publicationid: pubid: 979-8-3503-9042-1/24/\$31.00 ©2025 IEEE
@@ I Introduction
As power systems continue to integrate renewable energy resources, conventional Synchronous Generators (SGs) are progressively being replaced by Inverter-Based Resources (IBRs). This transition changes the dynamics of power systems, removing mechanical inertia and damping from the grid [1], which can jeopardize small-signal stability, defined as the ability of a power system to maintain synchronism under small disturbances [2].
@@
Several studies have examined how increased IBR penetration affects the electromechanical modes inherent to conventional power systems [3, 4, 5, 6]. Specifically, the authors of [3] and [4] discuss the impact that IBRs have on the existing modes in the system, highlighting that weak interconnections between IBRs and the grid, combined with decreasing system inertia and interactions between newly added and existing IBRs, can have a detrimental effect on system stability. Furthermore, [5] and [6] compared the converter-interface generation control strategies and examined their impact on grids with increasing penetration of IBRs, concluding that future power systems with a high share of non-synchronous generation may experience a reduction in small-signal stability, depending on the IBR control strategy.
@@
While most research has focused on the occurrence of oscillations and the role IBRs play in their emergence, comparatively few studies have explored the performance of existing Power System Stabilizers (PSSs) in grids with high IBR penetration, originally designed to mitigate electromechanical oscillations in SG-dominated systems [2]. The work in [3], [6], and [7] emphasized the need of re-tuning PSSs parameters since they were tuned decades ago for grids with dynamics dominated by SGs. Furthermore, the authors in [8] examined the transition from SG-dominated grids to IBR-dominated systems, including Grid Following (GFL) converters, highlighting potential instabilities arising from the interactions between GFL converters, PSSs, and Automatic Voltage Regulators (AVRs). Similar observations were made in [9], which identified PSSs and AVRs as potential sources of instability in systems with high IBR penetration.
@@
Using the Kundur Two-Area Test System as a benchmark case study, we investigate the effectiveness of legacy-tuned PSSs when a subset of the SGs are replaced by GFL converters. We use GFL inverter control in our case study as it is the prevailing control strategy today [10]. Subsequently, we investigate the efficacy of re-tuning the PSSs to restore system stability. Specifically, we investigate both “coordinated” and “local” PSS tuning approaches, in which communication between SGs is and is not allowed, respectively. We evaluate the performance of industry-standard, model-based PSS tuning methods (Residues and P-Vref) and identify additional requirements—specifically, coordination—needed for these methods to be effective.
Based on our tests on the benchmark Kundur Two-Area Test System, we draw the following findings:
  • The PSSs remaining in the system after replacing SGs with GFL converters require re-tuning, as their initial settings may lead to system instability.
  • When re-tuning just one PSS, the Residues method outperforms the P-Vref; however, the performance can still be unsatisfactory.
@@
  • When re-tuning both PSSs, the P-Vref outperforms the Residues method, which fails to stabilize the system when coordination is unavailable. Although re-tuning both PSSs with P-Vref stabilizes the test cases, it is not guaranteed to work for all systems.
@@
This paper is organized as follows: Section II provides an overview of small-signal stability and PSSs. Section III reviews the Residues and P-Vref methods. Section IV presents the case study and summarizes the main findings, while Section V concludes the paper.
@@ II-A System Definition
II-B Power System Stabilizer
To improve the small-signal stability of multi-machine power systems, it is common practice to integrate a PSS control loop into the generator’s AVR system, as shown in Fig. 1. The PSS adds $V_{\text{PSS}}$ to $V_{\text{error}}$ to compensate for $V_{\text{error}}$ oscillations and provides a damping-torque component in phase with the rotor speed deviation $\Delta\omega$. For more details on the PSS structure, we refer the reader to [2, Section 12.5].
@@ II-C Plant Transfer Function
Let us now consider a power system model defined by (2a) and (2b). We assume that a PSS is not connected and that the system has a single input $u$, corresponding to the AVR’s reference voltage $V_{\text{ref}}$, and a single output $z$, which represents the rotor speed $\omega$ of the analyzed SG, as depicted in Fig. 1. We transform the state vector $\mathbf{x}$ using the left eigenvector matrix $\mathbf{W=\begin{bmatrix}\mathbf{w}_{\mathrm{1}}^{\top}&\mathbf{w}_{\mathrm{2}}^{\top}&\dots&\mathbf{w}_{\mathrm{n}}^{\top}\end{bmatrix}^{\top}}$, where $\mathbf{w}_{i}\mathbf{A}=\lambda_{i}\mathbf{w}_{i},$ and obtain $\tilde{\mathbf{x}}=\mathbf{W}\mathbf{x}$. Substituting $\tilde{\mathbf{x}}$ into (2a) and (2b),
@@
$V_{\text{ref}}$++++++$V_{\text{error}}$ AVR Exciter Generator $E_{FD}$-- $K_{\text{PSS}}$ Gain $\frac{sT_{W}}{1+sT_{W}}$ High-pass Filter $\frac{1+sT_{1}}{1+sT_{2}}$ $\frac{1+sT_{3}}{1+sT_{4}}$ Lead/Lag Blocks$+$$V_{\text{PSS}}$$\min$$\max$LimiterPlant Transfer Function $G(s)$PSS Transfer Function $F(s)$$V_{C}$$\Delta\omega$
@@ III Methodology
Our analysis focuses on two tuning approaches: the Residues and P-Vref methods, both of which are model-based, as they require a power system model for tuning the PSSs. We refer the reader to [11, Section 6.2 and 5.8] for further details on the Residues and P-Vref methods, respectively.
@@ III-A Residues Method
[…]
@@
Using (7) and assuming the maximum phase compensation provided by each lead-lag block is approximately ±$\pm 60^{\circ}$, the required number of lead-lag blocks equals $N=1$ if $|\theta_{H_{\text{PSS}}}|<60^{\circ}$, $N=2$ if $60^{\circ}\leq|\theta_{H_{\text{PSS}}}|$ and $N=3$ if $120^{\circ}\leq|\theta_{H_{\text{PSS}}}|$. The parameters of each lead-lag block are then given as:
@@
Typically, the value of $T_{W}$ in the high-pass filter ranges from 1 to 20 s, depending on the target frequency range. Furthermore, the PSS gain $K_{\text{PSS}}$ may be determined by evaluating the damping provided to the mode $\lambda_{c}$ using the root locus of the TF (5).
@@ III-B P-Vref Method
The P-Vref Method uses a different TF, $G_{P\text{-}V_{\text{ref}}}(s)$, to tune the PSS parameters. This TF maps the reference voltage $V_{\text{ref}}$ to the generator’s active power output $P$ and is computed with the shaft dynamics disabled, i.e., $\Delta\omega=0$. After connecting $F(s)$, the PSS’s TF which maps $\Delta\omega$ to $V_{\text{PSS}}$, the compensated plant TF $G_{C}(s)$ can be written as [11, Section 5.8]
@@
The phase response of $G_{P\text{-}V\text{ref}}(s)$ may be obtained by running simulations in which the AVR’s reference voltage $V_{\text{ref}}$ is perturbed with a sinusoidal signal within the analyzed frequency range (0.1 - 10.0 Hz). The active power output of the corresponding SG is measured, and the phase shift between the input and output signals of $G_{P\text{-}V\text{ref}}(s)$ is calculated. This procedure is repeated for a number of frequencies within the desired range. During these simulations, the SG speed $\omega$ should remain constant, as deviations in $\omega$ affect $T_{e}$ and $P$. This can be achieved by setting the inertia constants $H$ of all SGs in the power system model to a sufficiently high value, thus ensuring that the shaft dynamics are disabled, i.e., $\Delta\omega\approx 0$.
The time constants of the lead/lag blocks $T_{1}$ to $T_{4}$ are tuned over the desired frequency range such that the phase response of the PSS best-matches the negative of the calculated phase response of $G_{P-V\text{ref}}(s)$. The high-pass filter $T_{W}$ is selected as done in the Residues method to target the frequencies of interest. After the time constants are selected, the PSS gain $K_{\text{PSS}}$ is tuned using root locus analysis of $W(s)$ from (5), as with the Residues Method.
@@ IV Results
Figure 2: Kundur 2-area system used in simulations: 0% IBR system (only SGs connected) and 50% IBR system (SG at Buses 1 and 3 replaced with GFLs).
@@
This section examines the influence of increasing IBR penetration on system stability and PSS tuning using dynamic simulations of the two-area system from [2] – a benchmark known for exhibiting both inter-area (0.1–0.7 Hz) and local (0.7–2.0 Hz) electromechanical oscillations – performed with the STEPSS simulation platform [12]. The system parameters can be found in [2, Example 12.6], corresponding to a thyristor exciter with high transient gain. The sixth-order SG model used in this study is from [13], while the speed governor represents a thermal turbine model with a reheater. The exciter and AVR model implemented across all generators is detailed in [14], and the PSS model is described in Section II-B. System loads are modeled as constant current for active power and constant impedance for reactive power.
@@
To assess system stability, we disturb the system with a 1% load increase at Bus 9 and observe the active power output trajectories of $G_{4}$ at Bus 4. Upon detecting instability, we re-tune the PSS at Bus 4 ($PSS_{4}$), identified by a participation factor and residue analysis as the most effective re-tuning location, without altering the remaining PSS at Bus 2 ($PSS_{2}$). If the system is still unstable after re-tuning $PSS_{4}$ (Section IV-B), $PSS_{2}$ is subsequently re-tuned.
@@ IV-B PSS Parameter Set B
The trajectories reveal that the system remains unstable after re-tuning $PSS_{4}$ using the P-Vref method, while the Residues-tuned $PSS_{4}$ exhibits poorly damped oscillations. The Residues-tuned system behavior can be attributed to the fact that the Residues method targets the unstable mode $\lambda_{1}$ (Fig. 4a) and shifts it to the left-half plane. However, as the gain increases, the $\lambda_{2}$ mode approaches instability. As a result, all stable cases for the selected gain $K_{PSS_{4}}$ exhibit a small damping ratio $\xi$. On the other hand, the P-Vref method cannot shift the unstable mode $\lambda_{1}$ to the left-half plane with any gain, as shown in Fig. 4b.
We conclude that, for Parameter Set B and tuning just $PSS_{4}$, the Residues method outperforms P-Vref, but still is unsatisfactory.
Figure 4: Movement of the critical poles in the root-locus plots for the initial PSS Parameter Set B under different re-tuning scenarios: (a) and (b) show the re-tuning of just $PSS_{4}$; (c) and (d) illustrate the sequential re-tuning of $PSS_{4}$ and $PSS_{2}$. The gray dotted lines indicate the damping ratio of the modes, defined as $\xi=-\frac{\sigma}{\sqrt{\sigma^{2}+\omega^{2}}}$. The stars indicate the open-loop pole positions (with $K_{\text{PSS}}=0$), whereas circles mark the zeros towards which the closed-loop poles migrate as the gain increases.1
@@
In uncoordinated PSS Re-tuning, each PSS re-tunes its PSS parameters simultaneously, without considering that the other PSS is also making adjustments. For Parameter Set B, we observed that uncoordinated Residues PSS Re-tuning did not stabilize the grid. This is due to the “wrong” A matrices used for tuning each PSS, which don’t consider that the parameters of the other PSS will change. Uncoordinated P-Vref, on the other hand, did stabilize the grid and produced similar behavior to the Sequential P-Vref method described next, and plotted in Fig. 3b with dashed red lines. The results are omitted here due to limited space.
@@
The dashed trajectories in Fig. 3b demonstrate that sequential tuning is effective with the P-Vref method, and does not change the performance of the Residues method significantly. The Residues method considers the full $A$ matrix, and thus is not well-suited for sequential tuning—in this test case, the retuning of $PSS_{4}$ tries to stabilize the entire system, without recognizing that $PSS_{2}$ will subsequently be retuned and can help. That is, the Residues tuning of $PSS_{4}$ establishes new dynamics for which $PSS_{2}$ has little impact. As shown by the root-locus for $PSS_{2}$ retuning in Fig. 4c, $PSS_{2}$ requires a large gain and still cannot push the pole far into the left-half plane.
@@ V Conclusion
11footnotetext: The root locus plots arise after the Residues or P-Vref method determines the time constants $T_{1}$ to $T_{4}$. The open-loop poles correspond to turning off the given PSS. In Fig. 4c, turning off $PSS_{2}$ destabilizes the system as the open-loop pole $\lambda_{4}$ is in the right-half plane. In Fig. 4d, turning off $PSS_{2}$ stabilizes the system as all open-loop poles are in the right-half plane.

### v2

@@ Abstract
publicationid: pubid: 979-8-3315-2503-3/25/\$31.00 ©2025 IEEE
@@ I Introduction
As power systems continue to integrate renewable energy resources, conventional Synchronous Generators (SGs) are progressively being replaced by Inverter-Based Resources (IBRs). This transition alters grid dynamics by removing mechanical inertia and damping [gfl], potentially jeopardizing small-signal stability, defined as the ability of a power system to maintain synchronism under small disturbances [Kundur1994].
@@
Several studies have examined how increased IBR penetration affects the electromechanical modes inherent to conventional power systems [cse_interaction, 9110789, epri_gfm, en15124273]. Specifically, the authors of [cse_interaction] and [9110789] discuss the impact that IBRs have on the existing modes in the system, highlighting that weak interconnections between IBRs and the grid, combined with decreasing system inertia and interactions between newly added and existing IBRs, can have a detrimental effect on system stability. Furthermore, [epri_gfm] and [en15124273] compared the converter-interface generation control strategies and examined their impact on grids with increasing penetration of IBRs, concluding that future power systems with a high share of non-synchronous generation may experience a reduction in small-signal stability, depending on the IBR control strategy.
@@
While most research has focused on the occurrence of oscillations and the role IBRs play in their emergence, comparatively few studies have explored the performance of existing Power System Stabilizers (PSSs) in grids with high IBR penetration, originally designed to mitigate electromechanical oscillations in SG-dominated systems [Kundur1994]. The work in [cse_interaction], [en15124273], and [beyond_low_inertia] emphasized the need of re-tuning PSSs parameters since they were tuned decades ago for grids with dynamics dominated by SGs. Furthermore, the authors in [uros_small_signal] examined the transition from SG-dominated grids to IBR-dominated systems, including Grid Following (GFL) converters, highlighting potential instabilities arising from the interactions between GFL converters, PSSs, and Automatic Voltage Regulators (AVRs). Similar observations were made in [stability_assesment], which identified PSSs and AVRs as potential sources of instability in systems with high IBR penetration.
@@
Using the Kundur Two-Area Test System as a benchmark case study, we investigate the effectiveness of legacy-tuned PSSs when a subset of the SGs are replaced by GFL converters. We use GFL inverter control in our case study as it is the prevailing control strategy today [dorfler_dominic_low_inertia]. Subsequently, we investigate the efficacy of re-tuning the PSSs to restore system stability. Specifically, we investigate both “coordinated” and “local” PSS tuning approaches, in which communication between SGs is and is not allowed, respectively. We evaluate the performance of industry-standard, model-based PSS tuning methods (Residues and P-Vref) and identify additional requirements—specifically, coordination—needed for these methods to be effective.
@@
Based on our tests on the benchmark Kundur Two-Area Test System, we draw the following findings. The PSSs remaining in the system after replacing SGs with GFL converters require re-tuning, as their initial settings may lead to system instability. When re-tuning just one PSS, the Residues method outperforms the P-Vref; however, the performance can still be unsatisfactory. In contrast, when both PSSs are re-tuned, the P-Vref method outperforms the Residues method, which fails to stabilize the system when coordination is unavailable. Although re-tuning both PSSs with P-Vref stabilizes the test cases, it is not guaranteed to work for all systems.
@@
The remainder of this paper is organized as follows. Section II provides an overview of small-signal stability and PSSs. Section III reviews the Residues and P-Vref methods. Section IV presents the case study and summarizes the main findings, while Section V concludes the paper.
@@ II-A System Definition
II-B Power System Stabilizers
To improve the small-signal stability of multi-machine power systems, it is common practice to integrate a PSS control loop into the generator’s AVR system, as shown in Fig. 1. The PSS adds $V_{\text{PSS}}$ to $V_{\text{error}}$ to compensate for $V_{\text{error}}$ oscillations and provides a damping-torque component in phase with the rotor speed deviation $\Delta\omega$. For more details on the PSS structure, we refer the reader to [Kundur1994, Section 12.5].
@@ II-C Plant Transfer Function
Let us now consider a power system model defined by (2a) and (2b). We assume that a PSS is not connected and that the system has a single input $u$, corresponding to the AVR’s reference voltage $V_{\text{ref}}$, and a single output $z$, which represents the rotor speed $\omega$ of the analyzed SG, as depicted in Fig. 1. We transform the state vector $\mathbf{x}$ using the left eigenvector matrix $\mathbf{W=\begin{bmatrix}\mathbf{w}_{\mathrm{1}}^{\top}&\mathbf{w}_{\mathrm{2}}^{\top}&\dots&\mathbf{w}_{\mathrm{n}}^{\top}\end{bmatrix}^{\top}}$, where $\mathbf{w}_{i}\mathbf{A}=\lambda_{i}\mathbf{w}_{i},$ and obtain $\tilde{\mathbf{x}}=\mathbf{W}\mathbf{x}$. Substituting $\tilde{\mathbf{x}}$ into (2a) and (2b) yields
@@
[…]
@@ III Methodology
Our analysis focuses on two tuning approaches: the Residues and P-Vref methods, both of which are model-based, as they require a power system model for tuning the PSSs. We refer the reader to [small_signal, Section 6.2 and 5.8] for further details on the Residues and P-Vref methods, respectively.
@@ III-A Residues Method
where $R_{c}=r_{c}\angle{\theta_{R_{c}}},\,\theta_{R_{c}}\in[0^{\circ},\,360^{\circ}]$. Assuming the PSS lead-lag blocks are equal, such that $T_{1}=T_{3}=T$ and $T_{2}=T_{4}=aT$, the maximum phase angle $\phi_{m}$ of a lead-lag block, which occurs at frequency $\omega_{m}=\frac{1}{T\sqrt{a}}$, equals
@@
Using (7) and assuming the maximum phase compensation provided by each lead-lag block is approximately $\pm 60^{\circ}$, the required number of lead-lag blocks equals $N=1$ if $|\theta_{H_{\text{PSS}}}|<60^{\circ}$, $N=2$ if $60^{\circ}\leq|\theta_{H_{\text{PSS}}}|$ and $N=3$ if $120^{\circ}\leq|\theta_{H_{\text{PSS}}}|$. The parameters of each lead-lag block are then given as:
@@
Typically, the value of $T_{\mathrm{W}}$ in the high-pass filter ranges from 1 to 20 s, depending on the target frequency range. Furthermore, the PSS gain $K_{\text{PSS}}$ may be determined by evaluating the damping provided to the mode $\lambda_{c}$ using the root locus of the TF (5).
@@ III-B P-Vref Method
The P-Vref Method uses a different TF, $G_{P\text{-}V_{\text{ref}}}(s)$, to tune the PSS parameters. This TF maps the reference voltage $V_{\text{ref}}$ to the generator’s active power output $P$ and is computed with the shaft dynamics disabled, i.e., $\Delta\omega=0$. After connecting $F(s)$, the PSS’s TF which maps $\Delta\omega$ to $V_{\text{PSS}}$, the compensated plant TF $G_{C}(s)$ can be written as [small_signal, Section 5.8]:
@@
The phase response of $G_{P\text{-}V\text{ref}}(s)$ may be obtained by running simulations in which the AVR’s reference voltage $V_{\text{ref}}$ is perturbed with a sinusoidal signal within the analyzed frequency range from 0.1 to 10.0 Hz. The active power output of the corresponding SG is measured, and the phase shift between the input and output signals of $G_{P\text{-}V\text{ref}}(s)$ is calculated. This procedure is repeated for a number of frequencies within the desired range. During these simulations, the SG speed $\omega$ should remain constant, as deviations in $\omega$ affect $T_{e}$ and $P$. This can be achieved by setting the inertia constants $H$ of all SGs in the power system model to a sufficiently high value, thus ensuring that the shaft dynamics are disabled, i.e., $\Delta\omega\approx 0$.
The time constants of the lead/lag blocks $T_{1}$ to $T_{4}$ are tuned over the desired frequency range such that the phase response of the PSS best-matches the negative of the calculated phase response of $G_{P-V\text{ref}}(s)$. The high-pass filter $T_{\mathrm{W}}$ is selected as done in the Residues method to target the frequencies of interest. After the time constants are selected, the PSS gain $K_{\text{PSS}}$ is tuned using root locus analysis of $W(s)$ from (5), as with the Residues Method.
@@ IV Results
Figure 2: Two-area system used in simulations: 0% IBR system (only SGs connected) and 50% IBR system (SG at Buses 1 and 3 replaced with GFLs).
@@
This section examines the influence of increasing IBR penetration on system stability and PSS tuning using dynamic simulations of the two-area system from [Kundur1994] – a benchmark known for exhibiting both inter-area (0.1–0.7 Hz) and local (0.7–2.0 Hz) electromechanical oscillations – performed with the STEPSS simulation platform [RAMSES]. The system parameters can be found in [Kundur1994, Example 12.6], corresponding to a thyristor exciter with high transient gain. The sixth-order SG model used in this study is from [STEPSS2023], while the speed governor represents a thermal turbine model with a reheater. The exciter and AVR model implemented across all generators is detailed in [VanCutsem2013], and the PSS model is described in Section II-B. System loads are modeled as constant current for active power and constant impedance for reactive power.
@@
To assess system stability, we disturb the system with a 1% load increase at Bus 9 and observe the active power output trajectories of SG $G_{4}$ at Bus 4. Upon detecting instability, we re-tune the PSS at Bus 4 ($PSS_{4}$), identified by a participation factor and residue analysis as the most effective re-tuning location due to its dominant involvement in the unstable mode, without altering the remaining PSS at Bus 2 ($PSS_{2}$). If the system is still unstable after re-tuning $PSS_{4}$ (Section IV-B), $PSS_{2}$ is subsequently re-tuned.
@@ IV-B PSS Parameter Set B
The trajectories reveal that the system remains unstable after re-tuning $PSS_{4}$ using the P-Vref method, while the Residues-tuned $PSS_{4}$ exhibits poorly damped oscillations. The Residues-tuned system behavior can be attributed to the fact that the Residues method targets the unstable mode $\lambda_{1}$ (Fig. 4a) and shifts it to the left-half plane. However, as the gain increases, the $\lambda_{2}$ mode approaches instability. As a result, all stable cases for the selected gain $K_{PSS_{4}}$ exhibit a small damping ratio $\xi$. On the other hand, the P-Vref method cannot shift the unstable mode $\lambda_{1}$ to the left-half plane with any gain, as shown in Fig. 4b, due to the high influence of $PSS_{2}$ on $\lambda_{1}$.
Based on this analysis, we conclude that for Parameter Set B and tuning just $PSS_{4}$, the Residues method outperforms P-Vref, but is still unsatisfactory.
Figure 4: Movement of the critical poles $\lambda_{j},j\in\{1,\dots,5\}$ in the root-locus plots for the initial PSS Parameter Set B under different re-tuning scenarios: (a) and (b) show the re-tuning of just $PSS_{4}$; (c) and (d) illustrate the sequential re-tuning of $PSS_{4}$ and $PSS_{2}$. The gray dotted lines indicate the damping ratio of the modes $\xi=-\frac{\sigma}{\sqrt{\sigma^{2}+\omega^{2}}}$. The stars indicate the open-loop pole positions (with $K_{\text{PSS}}=0$), whereas circles mark the zeros towards which the closed-loop poles migrate along the plotted trajectories as the gain increases1. The line styles correspond to those used in Fig. 3 for the respective PSS re-tuning scenario.
@@
In uncoordinated PSS Re-tuning, each PSS re-tunes its PSS parameters simultaneously, without considering that the other PSS is also making adjustments. For Parameter Set B, we observed that uncoordinated Residues PSS Re-tuning did not stabilize the grid. This is due to the “wrong” $\mathbf{A}$ matrices used for tuning each PSS, which don’t consider that the parameters of the other PSS will change. Uncoordinated P-Vref, on the other hand, did stabilize the grid and produced similar behavior to the Sequential P-Vref method described next, and plotted in Fig. 3b with dashed red lines. The results are omitted here due to limited space.
@@
The dashed trajectories in Fig. 3b demonstrate that sequential tuning is effective with the P-Vref method, and does not change the performance of the Residues method significantly. The Residues method considers the full $\mathbf{A}$ matrix, and thus is not well-suited for sequential tuning—in this test case, the retuning of $PSS_{4}$ tries to stabilize the entire system, without recognizing that $PSS_{2}$ will subsequently be retuned and can help. That is, the Residues tuning of $PSS_{4}$ establishes new dynamics for which $PSS_{2}$ has little impact. As shown by the root-locus for $PSS_{2}$ retuning in Fig. 4c, $PSS_{2}$ requires a large gain and still cannot push the pole far into the left-half plane.
@@ V Conclusion
11footnotetext: The root locus plots arise after the Residues or P-Vref method determines the time constants $T_{1}$ to $T_{4}$. The open-loop poles correspond to turning off the given PSS. In Fig. 4c, turning off $PSS_{2}$ destabilizes the system as the open-loop pole $\lambda_{4}$ is in the right-half plane. In Fig. 4d, turning off $PSS_{2}$ stabilizes the system as all open-loop poles are in the left-half plane.